OpenStudy (anonymous):

Combine like terms to simplify the expression: 7x + 3y – 2 + 6x – 1 + y2

OpenStudy (anonymous):

thats not an option

OpenStudy (anonymous):

7x + 6x = 13x 3y + 2y = 5y -2 + -1 = -3 13x + 5y - 3

OpenStudy (anonymous):

It's: 13x + y^2 + 3y - 3

OpenStudy (anonymous):

options are 2y3 + 13x – 3 3y3 + 13x2 – 3 y2 + 3y + 13x – 3 y2 + 3y + 13x2 – 3

OpenStudy (anonymous):

C is the same as my answer but ordered differently

OpenStudy (anonymous):

in other words the answer is C

OpenStudy (anonymous):

y2 + 3y + 13x – 3

OpenStudy (anonymous):

This website can help a lot: http://www.wolframalpha.com/input/?i=7x+%2B+3y+%E2%80%93+2+%2B+6x+%E2%80%93+1+%2B+y2+

OpenStudy (anonymous):

bbbbbbbb is correct it is C my bad i was wrong

OpenStudy (anonymous):

Type in your equation in the bar and it usually gives you whatever you need

OpenStudy (anonymous):

thanks so much guys!!!!!!